OpenStudy (anonymous):

MEDAL AND FAN: Choose the correct slope of the line that passes through the points (1, -3) and (3, -5).

OpenStudy (anonymous):

Answer choices: -1 1 0 1/2

Vocaloid (vocaloid):

slope = m = (y2-y1)/(x2-x1) let (1,-3) be (x1,y1) and let (3,-5) be (x2,y2)

OpenStudy (anonymous):

Can you explain further?

Vocaloid (vocaloid):

slope = m = (y2-y1)/(x2-x1) x1 = 1 x2 = 3 y1 = -3 y2 = -5

OpenStudy (hybrik):

So then -5-3/3-1 to get -8/2, simplify that to find your slope

Vocaloid (vocaloid):



OpenStudy (anonymous):

Thank you all :) I understand now

imqwerty (imqwerty):

If we r given any two points like- (x1,y1) and (x2,y2) then the slope m m=(y2-y1)/(x2-x1) Plug in x1, x2,, y1, y2 and get the answer

OpenStudy (sebaxtiangb):

u have to do rise over run